a charge of 0.000002C is taken from infinity to a point in an electric field,without changing its speed if work done by electric force is 0.00004J,then determine the potential at specified point. 

Grade:

2 Answers

anurag hundal
11 years ago

V=W/Q

V=0.00004/.000002

v=20!!!

Madhukar Thalore
11 years ago

hello,

work done = charge *potential

therefore, potential = 0.00004/0.000002=20Volt.